OpenStudy (anonymous):

Really need help here please. Solve the system by substitution. 2.5x-3y=-13 3.25x-y=-14

jagr2713 (jagr2713):

are there separate questions

OpenStudy (anonymous):

begin by solving for x or y in one of the equations, it will be in terms of the other variable

OpenStudy (dtan5457):

5x-3y=-13 25x-y=-14 I would take the second equation and make it -y=-14-25x which becomes y=14+25x Now just plug y=14+25x into the y value of 5x-3y=-13